### v1

# GAPS: Guiding Dynamic Android Analysis with Static Path Synthesis

Abstract.
Dynamically resolving method reachability in Android applications remains a critical and largely unsolved problem. Despite notable advancements in GUI testing and static call graph construction, current tools are insufficient for reliably driving execution toward specific target methods, especially those not embedded in a graphical component (e.g., libraries’ methods), a capability essential for tasks such as vulnerability validation, debugging, and behavioral analysis.
We present GAPS (Graph-based Automated Path Synthesizer), the first system that integrates static, method-guided call graph analysis with dynamic, interaction-driven execution. GAPS performs a lightweight backward traversal of the call graph, guided by data-flow analysis, to reconstruct paths reaching the target methods. These paths are then translated into instructions that guide runtime app exploration.
On the AndroTest benchmark, GAPS statically identifies paths to reach 88.24% of the target methods in just 4.27 seconds per app and dynamically reaches 57.44% of them. In contrast, state-of-the-art dynamic interaction tools show significantly lower reachability over three runs: APE, one of the best model-based GUI testers, achieves 12.82%, while GoalExplorer, a hybrid analysis tool, reaches 9.69%, and Guardian, an LLM-based UI automator, reaches 17.12%. Static analysis tools also fall short: FlowDroid and DroidReach identify paths to reach 58.81% and 9.48% of the targets, requiring 35.06 seconds and 23.46 seconds per app, respectively.
Finally, an evaluation on the 50 most downloaded real-world apps demonstrates GAPS’s practical utility in analyzing security-critical code under a realistic scenario. With an average static analysis time of 278.9 seconds, GAPS statically reconstructs paths to 62.03% of the target methods and dynamically reaches 59.86% of them.
1. Introduction
Android dynamic analysis is a process aimed at inspecting an Android application’s behavior during its runtime execution. It can be used to verify security vulnerabilities, measure performance, and detect malicious activities. Complementary to static analysis, which inspects the app without executing it, dynamic analysis involves running the app in a controlled environment, such as a real device, an emulator, or a sandbox, while monitoring its interactions with the system, network, and other components. Even though some apps, such as the Android Operating System (OS) services, can come without a Graphical User Interface (GUI), most apps require interaction with the available GUI components to change their status. Thus, dynamic analysis for Android apps mainly includes GUI testing tools. This choice already comes with an important limitation: such tools are able to reach only code within a GUI component and not any code inside an app. In fact, proving the reachability of an app target method is still an open research question (browserprivacy; monkeylimitations; androtest).
The evolution of Android GUI testing tools has gone through different phases, each one leveraging an incrementally sophisticated approach. The very first ones were pseudo-random input generators, such as Monkey (monkey), which lack of understanding of the app semantics and of the UI hierarchy, thus being inefficient for deep functional testing, but very useful for finding shallow random crashes in an app. Then, GUI testers started embedding models, such as finite state machines, to better reproduce common interaction scenarios and maximize exploration. Notably, APE (ape) is a model-guided GUI testing tool that has been recognized in prior studies for its effectiveness in improving code coverage and detecting GUI-related bugs (wang2021infrastructure; su2021benchmarking; akinotcho2024mobile). A hybrid solution, GoalExplorer (goalexplorer), builds a Screen Transition Graph (STG) to statically map relationships between activities and then guides automatic interaction using Stoat (stoat). The advent of machine learning and deep-learning techniques paved the way for GUI testers to autonomously discover new test paths following the guidance of optimization algorithms (ares). Finally, the introduction of Large Language Models (LLMs) allowed GUI testers to leverage semantic and contextual information to solve tasks closer to those performed by humans. Among recent advancements, Guardian (guardian) stands out for leveraging LLMs to guide UI exploration, showing competitive results across task completion benchmarks. Despite the multiple contributions in the field of Android automatic interaction (columbus; androtest; stochtesting; sapienz; timemachine; dynodroid), the efforts made so far have been mainly driven by either increasing the app code coverage or by replacing a real user in the completion of a task (e.g., “add an event to the calendar app”).
In this paper, we present our intuition that allows to finally guarantee method level reachability: the directed exploration of an app can be performed by following a path that connects an app entry point to the target code. Such path can be statically computed by traversing the app call graph, which shows the calling relationships between methods in the app.
At first, we evaluated the performance (i.e., target method reachability) of the best state-of-the-art tools designed for Android GUI interaction and call graph reconstruction. In particular, we randomly selected 50 target methods in each app among the 56 open-source ones [^1] https://github.com/samudoria/GAPS/tree/main/evaluation_data/androtest_results/ of the AndroTest (androtest) dataset, a standard benchmark in the field of Android automatic interaction, still used by current state-of-art (columbus; androtest; stochtesting; sapienz; timemachine; dynodroid; akinotcho2024mobile) . We, then, tested the performance of the selected tools. We started with the GUI testers, evaluating their ability to dynamically reach the target methods within a 5-minute timeout for each app, over three runs: on average, Guardian (guardian) was able to reach only 17.12% of the target methods, while APE (ape) and GoalExplorer (goalexplorer) achieved an average reachability rate of just 12.82% and 9.69% respectively, within the same time limit. In terms of call graph reconstruction, we chose FlowDroid (flowdroid), the best static analysis tool for such purpose (cg_soundness), and DroidReach (droidreach), the only available static path reconstruction tool: FlowDroid was able to reach 58.81% of the target methods, DroidReach was only able to reach 9.48% of the methods. Our findings confirm the lack of both static and dynamic solutions for reaching a target method in an app.
To this end, we design GAPS (Graph-based Automated Path Synthesizer), a framework that combines static analysis with guided dynamic execution to reach target methods in Android apps. Statically, GAPS performs target-oriented, context-sensitive call graph construction via backward traversal, rather than whole-program analysis, restricting its analysis to the subset of the program relevant to a given target method. It uses data-flow analysis techniques, such as points-to-analysis and constant propagation, to statically reconstruct partial execution paths using backward traversal from a target method, resolve conditional statements, and identify triggering UI elements. These results are compiled into high-level interaction plans that dynamically drive the app to execute the desired behavior.
Considering AndroTest, GAPS can generate at least one path for 88.24% of the target methods and automatically traverse such paths, reaching 57.44% of the target methods at runtime. Our static analysis module has an average completion time of 4.27 seconds, while our dynamic one never exceeded the 5 minutes timeout set.
Additionally, we evaluated GAPS on a real-world dataset consisting of the 50 most downloaded applications on the Google Play Store. This experiment serves as a use case scenario to demonstrate how GAPS can be employed in a realistic scenario to assess the reachability of security-relevant code. We began by running SPECK (speck), a rule-based static analyzer, to identify potential vulnerabilities and their corresponding method locations. These methods were then used as targets for GAPS to assess their reachability through both static and dynamic analyses. Our findings show that the static analysis module completed in an average of 278.9 seconds and was able to statically reach 62.03% of the target methods. The dynamic analysis achieved 59.86% of reachability.
The contributions of the paper are the following:
  • We design GAPS, a novel approach that integrates static and dynamic analyses to enable method-level reachability in Android apps. We release the framework and the evaluation data and setup in open source [^2] https://github.com/samudoria/GAPS/.
  • We evaluate GAPS against AndroTest, the state-of-the-art dataset for automatic interaction of Android apps. GAPS is able to generate at least one path to reach 88.24% of the target methods in 4.27 seconds on average, and to automatically reach 57.44% of them while never exceeding the 5-minute timeout.
  • We use GAPS in a real use case scenario, which is proving the reachability of potentially vulnerable methods in the 50 most downloaded real-world apps from the Google Play Store. With an average static analysis time of 278.9 seconds  we were able to statically reach 62.03% of the methods and achieve a runtime reachability of 59.86%.
2. Background
In this section, we describe the main methodologies used for automatically interacting with an Android app and for statically reconstructing its call graph.
2.1. GUI-Based Test Input Generators
Dynamic test input generators, often called GUI testers, explore Android apps by interacting with their GUIs. Their core functionality relies on repeatedly querying the device for the current view hierarchy, which is retrieved in XML format and lists all the interactable UI elements on the screen, such as buttons, checkboxes, text fields, and menus. Each GUI tester uses this structured representation to decide which element to interact with next. This decision process is guided by a strategy that may be pseudo-random (e.g., Monkey (monkey)), model-based (e.g., APE (ape)), machine-learning based (e.g., ARES (ares)), or semantic (e.g., Guardian (guardian)).
GUI testers are typically designed to maximize code coverage or detect UI-related bugs. Pseudo-random strategies rapidly issue events for lightweight testing, while model-based and machine-learning based tools interact heuristically with the UI, often guided by internal models of app states or by optimization algorithms, and prioritize breadth and novelty of interactions rather than reaching specific behaviors.
LLM-based tools follow a different paradigm. These systems are provided with a high-level task description (e.g., “save a note in the app”) and operate as agents that reason over the app’s interface to complete the task. Leveraging an LLM, they attempt to interpret screen layouts and prior interactions to plan and execute the sequence of actions.
Despite these advances, GUI testers, whether coverage-oriented or goal-driven, are fundamentally limited in their ability to consistently and specifically reach methods in an app, which results in an inefficient exploration. Moreover, they are definitely not able to reach methods that are not embedded in a GUI component (e.g., a library method).
2.2. Call Graph Generation
Call graphs are a core building block of static program analysis, since they capture the method invocation relationships within an app. In Android, call graph construction is especially challenging due to the platform’s event-driven architecture, components’ lifecycles, Inter-Component Communication (ICC), and dynamic behavior such as reflection.
Several tools, such as FlowDroid (flowdroid) and Amandroid (amandroid), attempt to build inter-procedural call graphs using context- and flow-sensitive analysis. These tools are computationally intensive and face scalability challenges due to their exhaustive exploration of all app edges. For instance, a previous study (mudflow) reported that even with 730GB of RAM and 24 hours of analysis time, FlowDroid failed to analyze certain applications. Additionally, Android call graphs are known to suffer from unsoundness, with missing or incorrect edges caused by unresolved dynamic behavior (cg_soundness). More specifically, many state-changing events for an Android app are not declared in its code, but are instead part of the Android framework, which implicitly invokes callback methods, such as onCreate, onClick (edgeminer; columbus). In addition, Android apps consist of multiple interacting components (i.e., Activities, Services, BroadcastReceivers), each governed by its own lifecycle. Accurately modeling these lifecycles and their transitions is critical for generating meaningful call graphs, but also difficult since they involve implicit invocations and asynchronous handlers to statically resolve. ICC introduces further complexity, as components interact by sending Intent messages that dynamically determine control-flow across components. Without proper modeling of ICC, the resulting call graph is incomplete and may miss valid transitions altogether (iccta). Finally, static analysis must also consider conditional logic. Even when a path exists between an entry point and a method, it may be guarded by a conditional statement.
Altogether, these challenges make it clear that full call graph construction for Android apps is not only resource-intensive, but also fundamentally limited in soundness and precision (cg_soundness). Moreover, complete call graph construction is not always necessary or even desirable when the analysis goal is to verify the reachability of a specific method. A full call graph introduces the risk of path explosion, especially in large apps with deeply nested lifecycles, asynchronous flows, and extensive use of libraries.
3. Related Work
In this section, we discuss current state-of-the-art research related to GUI testers, call graph generation and hybrid analysis.
3.1. GUI Testers
Automated GUI testing for Android has been approached with various techniques, but generally shares the goals of maximizing code coverage and finding UI bugs. Android’s built-in Monkey tool (monkey) is the archetype of random GUI testing. It simply issues pseudo-random UI and system events with minimal configuration, making it extremely easy to use. Variants like SmartMonkey (smartmonkey) and Dynodroid (dynodroid) are built on Monkey by adding simple heuristics to reduce wasted inputs. While random testers are scalable and easy to deploy, their lack of app-state awareness often leads to shallow coverage and repeated interactions with the same screens.
Model-based approaches construct an explicit GUI model (e.g., a finite state machine or transition graph) and systematically explore it. Notably, APE (ape) extends Monkey-like random testing with adaptive heuristics to improve coverage and efficiency. However, the models built at runtime are often incomplete or imprecise; they are unable to reason about internal logic and do not support ICC.
Reinforcement learning (RL) has been applied to adaptively guide GUI exploration. ARES (ares) and Q-Testing (qtesting) train models with rewards for novel screens and crashes, thus learning to prioritize unexplored transitions and incrementally improve coverage and bug discovery.
According to recent works (wang2021infrastructure; su2021benchmarking; akinotcho2024mobile), APE is recognized as one of the best-performing GUI testing tools. In subsection 4.2, we compared GAPS against APE, and our empirical evaluation shows that APE was able to reach only 12.82% of the target methods, in contrast to 57.44% achieved by GAPS.
3.2. LLM UI Automation
A rising trend in Android UI automation is to rely on LLMs to automatically interact with the device screen and complete a task specified by the user. GPTDroid (gptdroid) supplies the LLM with additional contextual information from the view hierarchy, along with a functionality-aware memory that maintains knowledge about past interactions. Autodroid (autodroid) integrates an initial random UI interaction to supply contextual information to the LLM and augment its performance. However, this approach is not exhaustive and scalable with bigger apps. VisionTasker (visiontasker) relies on a different approach by employing an object detection model to find UI elements from screenshots taken from the device. Although this strategy is more resistant to missing or misaligned elements, it is limited by what is currently shown on the screen and cannot properly process dynamic screens (e.g., games) or slow-loading apps. Guardian (guardian) has a replanning strategy that prevents invalid or repeated selection of GUI elements. This approach, coupled with state restoration, allows Guardian to favor app exploration. In subsection 4.2, we evaluated Guardian’s effectiveness in reaching targeted methods within AndroTest applications, achieving an average reachability of only 17.12% over three runs.
3.3. Call Graph Generation
FlowDroid (flowdroid) is a taint analysis tool that generates the app’s call graph to discover privacy and security leaks between defined sources and sinks. FlowDroid builds a call graph based on Soot (soot) and performs a flow- and context-sensitive analysis. However, FlowDroid did not model ICC in its initial version, leaving an essential contribution to call graphs undressed. This was later introduced with the integration of IccTA (iccta). Similarly, Amandroid (amandroid) is a taint analysis tool that builds an inter-procedural call graph and computes points-to information for all objects of an app component. FlowDroid and Amandroid employ data-flow analysis techniques to develop a call graph in a forward-visit approach, i.e., every branch is visited during construction. In addition, this approach has proven to be resource-intensive when integrated with the increasing complexity of apps (mudflow). GAPS aims at a more scalable approach: we build a partial inter-procedural call graph backward by starting from a target method invocation to limit the processing overhead.
DroidReach (droidreach) is the only known path reconstruction tool. Its path reconstruction techniques rely on the generation of a complete call graph using either the Androguard call graph creation tool (androguard_cg) or FlowDroid. In subsection 4.3, we compare GAPS against DroidReach and FlowDroid. In both cases, GAPS’s static analysis module proved to be faster and more effective at reaching target methods compared to its counterparts. Additionally, it provides specific app entry points, execution paths to satisfy conditional statements, and corresponding GUI events to guide automated interaction.
3.4. Hybrid analysis
Andrubis (andrubis) is a hybrid analysis tool created for detecting Android malware. Andrubis’ static analysis retrieves meta information (e.g., permissions, components, package name) and examines the Dalvik bytecode to extract a complete list of Java objects and methods. Subsequently, dynamic analysis is aided by information previously collected to interact with known entry points and measure code coverage, but it is not ultimately guided by the data collected during static analysis. AndroPyTool (andropytool) is another malware detection tool that combines static and dynamic analysis. It uses Androguard to extract essential features (e.g., package name, permissions, opcodes, API calls), FlowDroid to perform taint analysis, and samples are run inside a sandbox. However, dynamic analysis is not assisted by static analysis; instead, they are included as separate contributions in the output. GAPS aims to use static analysis information to provide path-sensitive instructions and guide automatic interaction. Evodroid (evodroid) proposes an evolutionary approach to the system testing of Android apps. It builds a Control Graph Model consisting of code segments used during GUI exercising to increase coverage. GAPS, instead, proposes a targeted approach to reach a specific method inside the app. Furthermore, Evodroid bases its static analysis on MoDisco (modisco), which requires the source code of the app to be provided, which is not always possible for most apps on the Google Play Store. GAPS can analyze any APK and DEX files as input and does not require the source code to operate. LetterBomb (letterbomb) employs a hybrid approach for automatically generating exploits for Android apps, specifically targeting Inter-Component Communication (ICC) vulnerabilities. The methodology combines a path-sensitive symbolic execution-based static analysis with software instrumentation and test oracles to identify and exploit weaknesses in the ICC interface. GAPS differs from LetterBomb in that it performs static analysis without relying on symbolic execution, which is known to be computationally expensive and often impractical for large, real-world apps. Instead, GAPS uses a lightweight static approach to reconstruct paths to a target method and then translates them into high-level instructions to guide UI exploration.
GoalExplorer (goalexplorer) is a hybrid system for automated, goal-driven exploration of Android applications. Its core idea is to first statically model the app’s UI screens and the transitions between them by combining results from modified versions of FlowDroid (flowdroid) and BackStage (backstage), creating an STG. This STG is then used to guide dynamic exploration, allowing the tool to navigate to a specific target of interest, such as an Android activity, an API call, or a program statement. In subsection 4.2, we have compared GoalExplorer against GAPS and it only reached 9.69% of the target methods on average over three distinct runs.
4. Motivation
To better understand the limitations of existing tools and evaluate the effectiveness of our approach, we structure our study around four research questions:
RQ1: How effective are GUI testers at reaching specific target methods?
RQ2: How effective is static analysis in reaching a method?
RQ3: How does GAPS perform compared to dynamic and static approaches?
RQ4: Can GAPS be applied to real-world apps and practical use cases?
4.1. Datasets & Setup
For the first three research questions (RQ1–RQ3), we used 56 applications from the AndroTest benchmark, for which we collected the apps’ latest version. These apps were instrumented using AndroLogs (samhi2024androlog) to measure method reachability by logging every time a method was executed. For each app, we randomly selected a consistent set of 50 target methods to be reached during testing, out of which only 34.39% of them are located inside an Activity.
For the final research question (RQ4), we selected the top 50 most downloaded apps from the Google Play Store.
Static analysis experiments are performed on a CloudLab (Duplyakin+:ATC19) Ubuntu 22.04 machine with 64 GB of RAM, while dynamic ones are performed on an x86-64 emulator running Android 13 and Pixel 2 with Android 11 when an ARM architecture is required.
4.2. RQ1: GUI testers effectiveness
To evaluate the effectiveness of GUI testers in reaching targeted methods within Android applications, we selected APE, recognized as one of the best-performing tools in recent studies (wang2021infrastructure; su2021benchmarking; akinotcho2024mobile), and Guardian, an LLM-based UI exploration tool that has outperformed other task-guided approaches (guardian). We also compared against GoalExplorer (goalexplorer), a hybrid analysis tool that statically builds an STG to guide the Stoat (stoat) stochastic GUI testing. The experiment was conducted using the AndroTest benchmark. Each app was executed for five minutes across three independent runs. We chose a 5-minute timeout for all competing tools, including GAPS, because we aimed to have an efficient dynamic analysis. Previous works (jordanfiveminutes; marvel) chose the same threshold. During execution, logs were collected to verify the dynamic reachability of the target methods. To execute Guardian, we tasked the LLM model to “interact with the application to maximize exploration”.
Considering the set of 50 target methods for each app, over three runs, GoalExplorer reached an average of 9.69% methods, APE 12.82%, and Guardian 17.12%. These results show how the three techniques differ in usability: APE can rapidly generate GUI events, but is unable to consistently build app states that can help discover more hidden parts of the app code. GoalExplorer uses the STG to guide Stoat by flagging activities when a target API is invoked. However, it faces limitations when the target methods reside in third-party libraries or when control-flow conditions are involved, as these often require more complex interactions beyond merely reaching the target activity. Finally, Guardian takes advantage of the LLM decision-making capabilities to systematically interact with UI elements in a goal-oriented manner, enabling it to uncover deeper execution paths. Despite this, its average reachability remains low, highlighting that semantic planning alone is insufficient without deeper control-flow awareness.
RQ1: GUI testers effectiveness Our evaluation highlights the limitations of dynamic interaction tools in reaching target methods. Both APE and Guardian show low success rates, with GoalExplorer reaching only 9.69%, APE 12.82%, and Guardian 17.12% of the predefined target methods across three runs. These findings highlight the need for a hybrid strategy that integrates dynamic execution with path-sensitive static guidance.
4.3. RQ2: Reachability with static analysis
We evaluated the effectiveness of static analysis tools in identifying paths to target methods using FlowDroid (flowdroid) and DroidReach (droidreach). FlowDroid is a well-established static analyzer that is actively maintained and was recently recognized as the best-performing tool for call graph generation (cg_soundness). DroidReach, on the other hand, is the only available tool that reconstructs paths by leveraging call graphs generated via AndroGuard (androguard_cg), applying graph-search algorithms to assess method reachability. In our setup, we improved DroidReach by restricting entry points to only exported components, whereas the original configuration considers all manifest-declared components as potential entry points. We note that DroidReach also supports FlowDroid-generated call graphs, but relies on an outdated version; thus, we excluded that option and ran the tools independently for comparison. We used FlowDroid in its CHA configuration, following a recent comparison study (cg_soundness). We also intended to include GoalExplorer in our static analysis comparison. However, GoalExplorer’s output is limited to the STG, which flags target activities, but does not report how many target APIs were actually found. In addition, the tool currently suffers from broken Maven dependencies, which prevented us from customizing it to produce the statistics required for a fair comparison. As a result, we had to rely on the pre-built executable released in 2024, which restricted our ability to adapt or extend the tool for our evaluation.
Experiments were conducted using the AndroTest benchmark. For each application, 50 randomly selected target methods were used. After execution, we examined the resulting call graphs and computed whether the target methods were reachable. Our results show that both tools face limitations. FlowDroid was able to reach only 58.81% of the target methods, while DroidReach performed worse, reaching just 9.48%. In terms of performance, FlowDroid had an average analysis time of 35.06 seconds, compared to 23.46 seconds for DroidReach.
Importantly, neither tool handles conditional logic along paths, meaning that reachability is assessed without validating whether the conditions required to execute those paths can actually be satisfied. Furthermore, the entry points used by both tools are either artificial or overly broad. More specifically, FlowDroid constructs a synthetic dummyMain method to simulate execution, but it provides no instructions for how to drive it or interpret paths from it. DroidReach, via AndroGuard, restricts itself to exported components declared in the manifest. Since Android applications can have entry points that are programmatically defined, this simplification might lead to missing relevant execution paths entirely.
RQ2: Reachability with static analysis Static analysis tools fail to identify a substantial portion of reachable target methods. FlowDroid reaches only 58.81%, and DroidReach performs even worse by reaching 9.48% of target methods. These findings show that current call graph generation tools are not effective for target-method reachability and motivate the need for more focused and goal-driven approaches.
5. Design
In this section, we first provide an illustrative example that we will use to explain how GAPS works. Specifically, we first describe an illustrative example to describe how GAPS’ static analysis pre-processing works and how its output guides the automatic interaction.
5.1. Illustrative Example
Here, we provide an example that will help describe each GAPS component. Furthermore, even with a simple example, it will be possible to immediately grasp the complexity of generating guidance for automatic interaction through static analysis. In our example, we want to build a path to reach the target_method in ActivityA from one of the app entry points, e.g., the MainActivity with its lifecycle. In 1, the MainActivity triggers the display of ActivityA (6 - 9) when the corresponding button is clicked (6). The button click triggers the creation of an Intent message (7) in the callback and dispatches it through the startActivity API (8). Furthermore, a BroadcastReceiver component is registered (5) and can be contacted by sending an Intent message by specifying the com.example.SET_ACCESS action.
1class MainActivity extends AppCompatActivity {
2 Button buttonMain = findViewById(button_main);
3
4 public void onStart() {
5 registerReceiver(new AccessReceiver(), new IntentFilter(‘‘com.example.SET_ACCESS’’));
6 buttonMain.setOnClickListener((view) -> {
7 Intent myIntent = new Intent(MainActivity.this, ActivityA.class);
8 startActivity(myIntent);
9 });
10 }
11}
Listing 1: The MainActivity starts ActivityA and registers AccessReceiver.
In 2, ActivityA invokes the target method (8) when a button is tapped (6), provided a conditional check in the onClick callback (7), involving the access variable, evaluates to true.
1class ActivityA extends AppCompatActivity {
2 Button buttonA = findViewById(button_a);
3 public static boolean access = False;
4
5 public void onStart() {
6 buttonA.setOnClickListener((view) -> {
7 if(access == true){
8 target_method();
9 }
10 });
11 }
12}
Listing 2: ActivityA contains the invocation to target_method.
Finally, in 3, the AccessReceiver class directly influences the reachability of the target method. Upon being contacted, it sets the access variable to true (3), thus affecting the required conditional statement.
1class AccessReceiver extends BroadcastReceiver {
2 public void onReceive(Context context, Intent intent) {
3 ActivityA.access = true;
4 }
5}
Listing 3: AccessReceiver sets the access variable to True.
5.2. Overview
The static phase in GAPS is designed to identify and reconstruct feasible execution paths that lead from application entry points to a specified target method, while also extracting contextual information necessary to guide subsequent dynamic interaction. This component operates with the goal of avoiding full call graph construction by focusing only on paths relevant to the target. To describe how GAPS statically analyzes applications, we use the illustrative example described in  subsection 5.1. The objective is to synthesize one or more execution plans that guide the app toward invoking target_method in ActivityA.
The first step is entry point resolution. GAPS identifies all externally invocable components by inspecting the application manifest and component declarations in an effort to map the app’s ICC. In this case, it recognizes both the MainActivity and the AccessReceiver component, the latter being contacted from outside the app via the broadcast action com.example.SET_ACCESS. The data-flow analysis behind this procedure is described in subsection 5.3.
Next, GAPS builds feasible paths between entry points and the target method using a lightweight, context-sensitive analysis. For each path, it performs backward slicing from the target method and tracks relevant control-flow transitions. More specifically, a partial call graph is built by concatenating method invocations, involving both explicit and implicit invocations. More details are provided in subsection 5.4. In this example, it reconstructs the sequence from target_method(), which is nested in a conditional branch and the click of the buttonA graphical element. Going further, ActivityA.onStart is only executed through the startActivity(Intent), further nested in the buttonMain click in MainActivity.onStart.
When a conditional statement is encountered, as in line 7 of 2, where access must be true, GAPS analyzes what influences its satisfiability. By performing data-flow analysis through a combination of constant propagation and points-to-analysis, it discovers that the static field ActivityA.access is set to true inside the onReceive method of AccessReceiver. Crucially, this receiver is registered in MainActivity.onStart, and it can be invoked externally by sending a broadcast intent. This process is further described in subsection 5.5. To ensure that the conditional check in ActivityA evaluates to true at runtime, GAPS includes a trigger for contacting the AccessReceiver as part of the execution plan. Furthermore, GUI analysis detects that both buttonMain and buttonA are connected to relevant method calls via click listeners, and therefore must be interacted with during execution. GAPS performs data-flow analysis to retrieve the GUI widgets’ identifiers. Specifically, it analyzes how the widget objects are initialized in the code, tracing their definitions to locate calls to findViewById. In our example, buttonMain and buttonA are both assigned through such calls, and the argument to each function (i.e., button_main and button_a) is recorded as the identifier of the corresponding UI element. Design choices around this process are explained in subsection 5.6.
Finally, GAPS synthesizes a high-level instruction sequence to guide the dynamic phase. This plan includes: (i) sending a broadcast Intent with action com.example.SET_ACCESS to invoke AccessReceiver, (ii) clicking GUI element with ID button_main in MainActivity to launch ActivityA, and (iii) clicking GUI element with ID button_a in ActivityA to reach target_method. The plan also tracks that the condition access == true must hold when the final interaction occurs.
5.3. Mapping Inter-Component Communication
GAPS employs a strategy to map ICC based on the understanding that sending an Intent object implies passing it as an argument to a specific Android API call. The primary objective is to extract information regarding both the sender and receiver of the message exchanged through ICC. This way, components that are reachable can be used as app entry points during dynamic interaction, either explicitly (via a fully qualified class name) or implicitly (through an action field).
As shown in Algorithm 1, the initial step undertaken by GAPS involves the static retrieval of all ICC information attributed to the components.
Input: components: Declared components from the manifest and code
Output: icc: Mapping of components to intent actions, class names, and paths
1
2icc $\leftarrow$ empty dictionary
3
4foreach $comp\in components$ do
5 if is_exported(comp)is\_exported(comp) then
6      icc[comp] $\leftarrow$ comp.name
7
8   end if
9 if has_intent_filter(comp)has\_intent\_filter(comp) then
10      icc[comp] $\leftarrow$ comp.intent_actions
11
12   end if
13
14 end foreach
15
16dynamic_receivers $\leftarrow$ find_dynamic_receivers()
17
18foreach $recv\in dynamic\_receivers$ do
19   reg_paths $\leftarrow$ get_registration_paths(recv)
20   info $\leftarrow$ data_flow_analyze(recv, reg_paths)
21 if infoinfo contains action then
22      icc[recv.class] $\leftarrow$ info.action
23
24   end if
25
26 end foreach
27
28icc_paths $\leftarrow$ find_icc_related_paths()
29
30foreach $path\in icc\_paths$ do
31   intent $\leftarrow$ analyze_intent(path)
32   target $\leftarrow$ data_flow_resolve(intent)
33 if targettarget is not null then
34      icc[target] $\leftarrow$ path
35
36   end if
37
38 end foreach
39
return iccicc
Algorithm 1 ICC Mapping
Initially, the input components are inspected to determine whether they are exported or associated with intent filters (2-9), allowing GAPS to identify valid application entry points beyond the main activity. Next, the system analyzes dynamically registered BroadcastReceivers (10), such as the one defined in 3, which can be programmatically created and thus require data-flow analysis to be mapped. The function find_dynamic_receivers (10-12) first locates code sections that register these components, and then applies data-flow analysis (13) to extract the relevant metadata–specifically, the class names and associated IntentFilter objects. An IntentFilter defines invocation parameters such as the action, which is critical for triggering the corresponding receiver (15). This information is used to populate the icc dictionary, which maps destination components to their respective action values or fully qualified class name. To comprehensively model all dynamically reachable components, GAPS proceeds to analyze paths involving inter-component communication (19-25). This involves inspecting the construction of Intent objects (20) and applying data-flow analysis to resolve their targets (21). These resolved paths are incorporated into the same icc data structure (23), enabling GAPS to link ICC operations to execution flows.
In the illustrative example, the MainActivity uses the registerReceiver method to instantiate a broadcast receiver, that will accept intent with the “com.example.SET_ACCESS” action, as defined by the corresponding IntentFilter. Additionally, the startActivity API allows to start the ActivityA.
5.4. Backward Generation of the Call Graph
GAPS adopts a goal-directed, demand-driven approach to call graph construction that departs from traditional whole-program analyses, such as the one used by Flowdroid. Instead of building a complete call graph upfront, GAPS performs a targeted, backward traversal starting from a specified method of interest. This traversal is both inter-procedural and context-sensitive, collecting only those method calls that lie along feasible paths leading to the target. The resulting structure is a partial call graph slice that is sufficient for reachability reasoning but avoids the scalability bottlenecks of global analyses. GAPS also incorporates data-flow analysis to capture branch conditions and control dependencies essential for reconstructing realistic execution paths, as well as actionable entry points and GUI event identifiers. The pseudocode for the partial call graph generation can be consulted in Algorithm 2.
The process starts from the target method and iteratively traverses backwards through its potential callers. This is implemented using a queue of nodes to process (2 and 3), where the current node being processed is removed from a queue (6) and all its statically discoverable callers are identified via find_new_nodes (10). These new nodes are added to the call graph and appended to the queue (11). When an entry point is encountered, it serves to terminate the backward traversal, and no further predecessor nodes are added. Then, nodes that are processed are added to the seen_nodes set (12), to avoid reprocessing and ensure termination (7). This backward expansion continues until no new nodes are found. At that point, the constructed call graph contains a slice of the original application call graph that is specifically focused on reaching the target. Once the partial call graph is complete, GAPS invokes a depth-first traversal (dfs_visit) (14) to extract all feasible paths from known application entry points to the target method. These paths are further enriched by resolving conditional branches (16 and 17) and detecting associated GUI elements (18), which are necessary to trigger the GUI events.
Input: target_method: Target method
Output: result_paths: Set of paths inter-procedurally reconstructed
1
2result_paths $\leftarrow$ set()
3 nodes_queue $\leftarrow$ queue()
4 nodes_queue.add(target_method)
5 seen_nodes $\leftarrow$ set()
6 cg $\leftarrow$ graph()
7
8foreach $current\_node\in nodes\_queue$ do
9 if $current\_node\in seen\_nodes$ then
10    continue
11
12   end if
13  new_nodes $\leftarrow$ find_new_nodes(current_nodecurrent\_node)
14   current_node $\leftarrow$ add_new_nodes(new_nodesnew\_nodes, cgcg, nodes_queuenodes\_queue)
15   seen_nodes.add(current_nodecurrent\_node)
16
17 end foreach
18
19result_paths $\leftarrow$ dfs_visit(cgcg, target_methodtarget\_method)
20
21foreach $path\in result\_paths$ do
22   conditional_paths $\leftarrow$ find_conditional_paths(pathpath)
23   path $\leftarrow$ merge_paths(pathpath, conditional_pathsconditional\_paths)
24   path $\leftarrow$ find_GUI_elements(pathpath)
25
26 end foreach
27
return result_pathsresult\_paths
Algorithm 2 Partial call graph generation
5.5. Linking paths that satisfy conditional statements
During path reconstruction, as mentioned in subsection 5.4, it is possible to analyze the conditional statements and retrieve paths that satisfy them. Currently, GAPS supports conditional statements that involve: (i) variables, primitive types (e.g., int, String, float), (ii) objects, which can be compared against null, (iii) methods by considering their return values. GAPS combines point-to-analysis and constant-propagation to retrieve the parties concerned and the constant values they can assume. More specifically, points-to analysis is a static program analysis technique used to determine the set of memory locations (objects or variables) that a pointer or reference variable may refer to, while constant propagation is a compiler optimization technique that aims to replace variables with their constant values whenever possible. Algorithm 3 illustrates how they are used to retrieve paths that satisfy conditional statements.
Input: cond: The conditional statement to analyze
path: The execution path containing the conditional statement
Output: branch_paths: List of paths annotated with the conditional outcome
1
2lhs, rhs $\leftarrow$ extract_operands(cond)
3
4lhs_points $\leftarrow$ points_to_analysis(lhs, path)
5 rhs_points $\leftarrow$ points_to_analysis(rhs, path)
6
7lhs_vals, lhs_paths $\leftarrow$ constant_propagation(lhs_points)
8 rhs_vals, rhs_paths $\leftarrow$ constant_propagation(rhs_points)
9
10branch_paths $\leftarrow$ empty list
11
12foreach $val\_l,p\_l\in(lhs\_vals,lhs\_paths)$ do
13 foreach $val\_r,p\_r\in(rhs\_vals,rhs\_paths)$ do
14    if verify_condition(val_l,val_r,cond)verify\_condition(val\_l,val\_r,cond) then
15         merged_path $\leftarrow$ merge_paths(p_l, p_r, path)
16         branch_paths.add(merged_path)
17
18      end if
19
20   end foreach
21
22 end foreach
return branch_pathsbranch\_paths
Algorithm 3 Conditional Statement Resolution
Initially, the conditional statement and its containing path are supplied in input. The algorithm begins by extracting the left and right operands (lhs and rhs, respectively) from the conditional expression (1). For each operand, it performs points-to analysis to trace where the variable or reference may have originated in the program (2 and 3). These origin points are then used for constant propagation, which computes all possible constant values the operands may hold, along with the paths where these values are assigned (4 and 5). Once the constant values and their corresponding assignment paths are obtained, the algorithm performs a pairwise comparison of all value combinations (7-13). For each combination, it merges the two associated assignment paths with the current path context and verifies whether the condition could evaluate to true (9). If the condition is satisfiable under that value pair, the merged path is added to the output list of feasible, conditionally valid paths (10-11). The resulting branch_paths captures all statically inferred control-flow paths that satisfy the given conditional statement.
The illustrative example contains a conditional statement involving the access variable that will be analyzed to find the paths that set its value to True. This happens by contacting the AccessReceiver component. A separate path will be reconstructed and prepended to the main one to reproduce this.
5.6. GUI Event Retrieval
To support guided dynamic execution, GAPS must identify the specific GUI elements responsible for triggering event handlers within the app. This enables the generation of interaction plans that simulate user actions accurately. The retrieval of these GUI elements is performed statically and relies on a layered application of points-to analysis. Algorithm 4 illustrates the procedure.
The process begins by detecting GUI event handlers, such as those associated with callbacks like onClick() or onItemSelected(), that happen in a recognized activity. GAPS scans the paths for listener registrations, including explicit calls to methods like setOnClickListener(). Once a handler is found, points-to analysis is used to resolve the Android object instance linked to the callback (2), such as a Button, MenuItem, or other view element. After identifying the GUI object, the paths where it is involved are retrieved (3), and an iterative round of points-to analysis is used to trace its instantiation back to common methods such as findViewById(). From this trace, GAPS extracts the associated resource identifier (e.g., button_main) by analyzing the arguments passed to the call (5). The final outcome of this analysis is a list mapping the activity and the view identifier, representing a GUI event. These events are later embedded into high-level interaction instructions, allowing the dynamic executor to match statically derived GUI elements with actual runtime views.
Input: handler: GUI event handler method
activity: Activity where GUI event occurs
path: Execution path containing the event registration
Output: gui_bindings: Mapping from activity to resource ID
1
2gui_bindings $\leftarrow$ empty list
3
4gui_obj $\leftarrow$ points_to_analysis(handler, path)
5
6gui_obj_paths $\leftarrow$ find_paths(gui_obj)
7
8foreach $path\in gui\_obj\_paths$ do
9   gui_id $\leftarrow$ points_to_analysis(path)
10   gui_bindings.append(activty, gui_id)
11
12 end foreach
13
return gui_bindingsgui\_bindings
Algorithm 4 GUI Event Resolution
In the illustrative example, this process allows for the retrieval of the required interactions along the paths. More specifically, the GUI event related to ID button_main in MainActivity to launch ActivityA, and the one in ActivityA related to ID button_a.
6. Implementation
In this section, we provide more details about GAPS’ implementation: the call graph generation and path reconstruction, Graphical Elements’ ID retrieval, and the automatic interaction module.
6.1. Call Graph Generation and Path Reconstruction
At its core, GAPS uses the Androguard (androguard) Python library to load and analyze the APK or DEX files in input through the AnalyzeAPK and AnalyzeDex (analyzeapk) built-in functions. Then, GAPS analyzes the MethodAnalysis construct for each method to access its basic blocks. Every basic block obtained is represented by a sequence of smali instructions, an intermediate representation of the app bytecode. Internally, GAPS uses the networkx (networkx) Python library to create and model the graph. Subsequently, we use the all_shortest_paths (all_shortest_paths) method from the networkx library to retrieve the paths connecting the source nodes and the entry points. Since solving implicit flows is still considered an open research problem (cg_soundness) that falls out of the scope of this paper, we have added callback mappings provided by popular works in Android static analyzers, such as EdgeMiner (edgeminer) and Soot (soot) through its virtual edges (virtualedges). In this way, we can find the corresponding constructor or method initializer and partially solve implicit flows. Data-flow analysis is performed following the smali code representation, which provides a rigorous register-based syntax that we modeled to adapt the points-to-analysis and constant propagation. For example, to track the transition to a new activity, when an Intent is constructed, the last argument passed to the constructor corresponds to the destination activity, which matches the last register in smali syntax. This data-flow-aware analysis is also essential for linking paths that satisfy conditional statements. Moreover, through constant propagation, a conditional statement involving primitive-type variables, objects, or method return values is translated to compare constant values to test its satisfiability.
6.2. Static Output and Dynamic Guidance
GAPS uses Apktool (apktool) to decompile apps and retrieve GUI resource IDs during static analysis. To convert them from the smali numerical representation, GAPS consults a file generated by Apktool called public.xml, where they can be converted into a literal one. GAPS can generate high-level instructions in JSON files, consisting of actionable entry points and sequences of graphical element IDs to interact with, along with the Activity name where they reside. Additionally, a call sequence is also included. These results can be used in two ways. First, they can serve as the standalone output of the static analysis phase: for each target method, GAPS lists all the paths found during analysis, together with their reconstructed call sequence and the corresponding high-level interaction plan. Second, they can directly guide dynamic analysis. In this case, each plan begins with instructions that can be executed through adb to send an Intent and trigger the correct entry point discovered during the static analysis. The plan then continues with a sequence of UI element IDs and the Activity in which they appear, specifying the concrete actions to perform on the app interface in order to reach the target method. When conditional statements are resolved during the static phase, the relative call sequences and their associated high-level instructions are prepended to the plan, ensuring that the conditions are satisfied before continuing with the UI interactions.
6.3. Automatic Interaction
The high-level instructions generated through the static analysis allow for dynamic traversal of the paths by (i) using one of the app’s entry points and (ii) graphically interacting with the GUI. More specifically, the GAPS dynamic module relies on the AndroidViewClient (androidviewclient) Python library to find the graphical elements through the findViewById (avc1) and touch (avc2) APIs, given their literal identifiers, and subsequently interact with them. Additionally, Frida (frida), a dynamic instrumentation framework, can be used to apply a hook on the target method and get a confirmation when it is reached.
The previous approach comes with some limitations for apps that display unpredictable elements in an overlay on the screen. For example, many apps can show dynamically rendered content or require the supply of textual input. To address the unpredictability of some apps’ UI components, we integrated Guardian (guardian) in the GAPS’ automatic interaction module: when GAPS cannot locate an activity or widget on the screen, Guardian guides the interaction with the app until both the widget and the activity appear, enabling GAPS to resume its automatic interaction under path guidance. This choice is motivated by the comparison in subsection 4.2, where Guardian performed better than APE and GoalExplorer thanks to its LLM capabilities, which enable goal-driven interaction planning and more effective UI navigation, including the generation of textual input. For instance, applications that require semantic planning, such as Yahtzee, where the number of rounds and players must be specified in the very first activity, are better suited for exploration through LLM-based UI interaction tools like Guardian. We include Guardian as an optional component, which can be disabled or replaced with more advanced tools as they become available in the future. Furthermore, GAPS uses memorization to save the sequence of GUI events performed by Guardian when it is able to transition between two activities, allowing GAPS to reproduce the steps independently.
7. Results
In this section, we present the results of our experimental evaluation of GAPS. In subsection 7.1, we compare GAPS against representative static and dynamic tools using the AndroTest benchmark, evaluating both its static path synthesis and dynamic guided execution capabilities. Finally, in subsection 7.2 we present a case study on 25 real-world applications attempting to reach security-relevant code. The target methods, in this case, are provided by the SPECK (speck) rule-based static analyzer.
7.1. RQ3: Effectiveness of GAPS Compared to Existing Tools
To evaluate whether GAPS improves the reachability of target methods over state-of-the-art tools, we compare it against two static analyzers (FlowDroid and DroidReach) and three dynamic GUI testers (GoalExplorer, APE, and Guardian). GAPS is evaluated in both static and dynamic configurations: the static phase measures the percentage of methods for which GAPS can generate paths, while the dynamic phase measures actual runtime reachability during guided execution.
Table 1. Comparison of target-method reachability and analysis time across static and dynamic tools.
Tool Reachability (%) Avg. Analysis Time
Static Analysis Tools
DroidReach 9.48% 23.46 seconds
FlowDroid 58.81% 35.06 seconds
GAPS (static) 88.24% 4.27 seconds
Dynamic Interaction Tools
GoalExplorer 9.69% 00:05:00 (Timeout)
APE 12.82% 00:05:00 (Timeout)
Guardian 17.12% 00:05:00 (Timeout)
GAPS (dynamic) 57.44% 00:03:15
Table 1 presents a comparative analysis of target-method reachability and average analysis time for static and dynamic tools. GAPS significantly outperforms both classes of tools. Its static pre-processing achieves 88.24% reachability by generating partial paths that satisfy known conditions, entry points, and provides GUI events. Furthermore, its analysis time, on average, is 81.8% and 87.8% faster compared to DroidReach and FlowDroid, respectively. Meanwhile, its dynamic component, guided by statically synthesized interaction plans, reaches 57.44% of targets during execution, substantially higher than both APE and Guardian.
On average, Guardian was invoked four times per app, which corresponds to contributing to reaching 9.3% of the methods, indicating that GAPS primarily relies on its static analysis module for guidance, with Guardian used only as a fallback.
RQ3: Effectiveness of GAPS Compared to Existing Tools GAPS outperforms static and dynamic state-of-the-art tools both in terms of analysis time and target-method reachability. Its static phase covers 88.24% of target methods with an average analysis time of 4.27 seconds, while dynamic execution reaches 57.44%.
7.2. RQ4: GAPS’ Applicability to Real-World Use Cases
To demonstrate the practical utility of GAPS, we evaluated its performance on the 50 most downloaded real-world applications from the Google Play Store, collected in August 2025. The full list of applications and the results can be found in our repository [^3] https://github.com/samudoria/GAPS/tree/main/evaluation_data/real_world_apps_results/. It should be noted that, in this case, AndroLogs (samhi2024androlog) failed to repackage the real-world apps to add code coverage feedback. Instead, we used GAPS’ Frida integration to hook the target methods and log whenever they were reached.
GAPS achieved a static reachability of 62.03% of the target methods, with an average analysis time of 278.9 seconds (less than 5 minutes) per app. This performance highlights GAPS’s ability to scale to large, complex applications while maintaining efficiency. The lower reachability compared to the AndroTest benchmark can be attributed to the increased complexity and obfuscation techniques employed in commercial apps, as well as the use of third-party frameworks which obscure traditional control-flow patterns. During dynamic execution, GAPS successfully reached 59.86% of the target methods. The gap between static and dynamic reachability arises from challenges such as when a complex app state is required (e.g., performing an app payment), dynamic UI rendering, and the need for specific input data.
RQ4: GAPS’ Applicability to Real-World Use Cases GAPS was evaluated on 50 top Google Play Store apps, demonstrating practical utility and scalability in real-world scenarios. The tool achieved 62.03% static reachability of target methods with an average analysis time of 278.9 seconds (less than 5 minutes per app), and 59.86% dynamic reachability.
7.3. Discussion and Limitations
While GAPS demonstrates significant improvements in targeted method reachability, it comes with some limitations.
First, its reliance on static bytecode analysis restricts its effectiveness for apps built with frameworks like Flutter or React Native, where the app’s logic resides outside traditional Dalvik code. Furthermore, apps with an obfuscated call graph can lead to path explosion and significantly slow down analysis time, which can be noted from the higher analysis time of real-world apps. Another obstacle to path reconstruction is dead code located in libraries, which can worsen GAPS’ performance. Finally, call graph generation is known to suffer from unsoundness (cg_soundness), caused by untracked implicit flows. GAPS uses the same callback mappings used by FlowDroid (flowdroid) and Soot (soot), and also includes those generated by EdgeMiner (edgeminer), but might still suffer from unsoundness when an unexpected implicit flow is met. The analysis of the Android Framework to mine implicit edges for static analysis tools is an orthogonal problem that is out of scope for this work. Our current approach does not support Jetpack Compose. Unlike XML-based layouts, Jetpack Compose defines UI elements directly in Kotlin code and attaches event handlers through lambdas rather than OnClickListener objects. Furthermore, UI state can change dynamically through recomposition. These differences make data-flow analysis for Jetpack Compose more complex. Extending static analysis to reliably extract GUI elements and events from Jetpack Compose remains an open research problem, and we consider this an important direction for future work.
In some cases, we cannot dynamically attest the reachability. Complex interactions might be required to fully traverse a path: Hot Death and Yahtzee are games where some of the randomly selected target methods can only be reached upon winning a game. Finally, most apps have paths with intent filters as an entry point, which can have permissions or be owned by the system (e.g., BOOT_COMPLETED), or could require data to be passed along with the Intent. Retrieving the required data parameters and automatically generating Intents is an orthogonal problem to dynamic analysis that falls out of the scope of this paper.
Despite these challenges, GAPS’s hybrid approach balances precision and scalability, offering a practical foundation for future enhancements.
8. Conclusions
In this paper, we present GAPS, a tool aimed at path reconstruction through the generation of Android apps’ call graph, that also offers a module to traverse the identified paths automatically. In our evaluation, we show that GAPS can reconstruct at least one path for 88.24% target methods and traverse 57.44% of such paths for apps belonging to the AndroTest benchmark with an average analysis time of 4.27 seconds. In contrast, state-of-the-art tools achieve significantly lower reachability: GoalExplorer (9.69%), APE (12.82%), Guardian (17.12%), FlowDroid (58.81%), DroidReach (9.48%), with longer or less targeted analysis phases. On the other hand, the results on the real-world apps dataset report that GAPS can reconstruct paths for 62.03% of the target methods, averaging an analysis time of 278.9 seconds per app, and dynamically proving the reachability for 59.86% of them.
Overall, we can conclude that GAPS can fill the gap between static and dynamic analyses, guiding a targeted app exploration and solving the method reachability problem.
9. Data Availability
We release GAPS’ source code, its evaluation data, and setup in open-source  [^4] https://github.com/samudoria/GAPS/. Additionally, we have included instructions to reproduce GAPS’ results on the AndroTest benchmark.